determine the shortest cable ABC that can be used to lift the 910lb beam if the maximum force the cable can sustain is 1500lbs

10 ft
Answer #1

here,

the maximum tension in the rope , t = 1500 lbs

weight of the beam , w = 910 lb

let the angle be theta

2 * t * sin(theta) - w = 0

2 * 1500 * sin(theta) - 910 = 0

theta = 17.66 degree

length of the cable ABC , l = 2 * 5 * cos(theta)

l = 2 * 5 / cos(17.66)

l = 10.49 m

the length of cord ABC is 10.49 m